OpenStudy (anonymous):

f(x)=(4x^2)/(x^2-1) Find the points of intersection, if any, of the graphs of y = f (x) and y = 4.

sam (.sam.):

You got equations y = f (x) y = (4x^2)/(x^2-1) and y = 4 you can solve them simultaneously

OpenStudy (anonymous):

thank you. but i can't solve this

sam (.sam.):

Try graphing ,

OpenStudy (agent0smith):

Set them equal to each other \[\large 4 = \frac {(4x^2)}{(x^2-1)}\] Try solving for x, if possible.

OpenStudy (agent0smith):

Try multiplying both sides by (x^2-1)

sam (.sam.):

Seemed to be no solution or no intersection

sam (.sam.):

Here's the graph

OpenStudy (agent0smith):

\[\large 4(x^2-1) =4x^2\]

OpenStudy (anonymous):

become 4x^2-4=4x^2 -4x^2+4x^2-4=0 then -4=0

OpenStudy (agent0smith):

Correct, which is not true, so no solution.

OpenStudy (anonymous):

i soleve taht . but i ask not another solve

OpenStudy (anonymous):

thank you sam and agent0smaith